# Auto-Vocabulary Semantic Segmentation

### v2

@@ Abstract
Open-ended image understanding tasks gained significant attention from the research community, particularly with the emergence of Vision-Language Models. Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, they operate without the need for training or fine-tuning. However, OVS methods typically require users to specify the vocabulary based on the task or dataset at hand. In this paper, we introduce Auto-Vocabulary Semantic Segmentation (AVS), advancing open-ended image understanding by eliminating the necessity to predefine object categories for segmentation. Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using enhanced BLIP embeddings, which are utilized for segmentation afterwards. Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated class names and their corresponding segments. Our method sets new benchmarks on datasets such as PASCAL VOC and Context, ADE20K, and Cityscapes for AVS and showcases competitive performance to OVS methods that require specified class names. All code is publicly available here.
Keywords:
Auto-Vocabulary Semantic Segmentation Open-Vocabulary Semantic Segmentation Semantic Segmentation
@@ 1 Introduction
Semantic segmentation is a computer vision task that involves grouping pixels into coherent regions based on their semantic categories. While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes. In contrast, humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories. Recent studies have focused on the development of segmentation models to meet similar capabilities [24, 20], with models that leverage Vision-Language Models (VLMs) as an emerging category [15, 10, 7, 17, 5, 40, 3, 37, 39]. Such VLMs learn rich multi-modal features from large numbers of image-text pairs.
@@
Figure 1: Different Semantic Segmentation Tasks in Comparison. In traditional Semantic Segmentation, an image is segmented into fixed, predefined dataset classes used during annotation. In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented. In Auto-Vocabulary Segmentation, relevant object categories are generated from the image automatically, enabling a true open-ended understanding of scenes without a human in the loop.
Figure 2: AutoSeg Exemplary Results. While established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment object categories beyond the fixed-set ground truth, such as cub, dachshund or pagoda. Images are from the PASCAL VOC [9] and ADE20K [41] datasets.
Given the scale, training VLMs comes at a extremely high computational cost. Because of this, it is common to build upon pre-trained VLMs such as CLIP [25] or BLIP [16]. However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image. Previous attempts have employed a two-stage approach where masked regions are fed to the VLMs to produce CLIP embeddings, which are subsequently correlated with text encodings of text prompts to perform classification [38, 17, 7]. Other works directly align per-pixel embeddings with text embeddings retrieved from CLIP [15] based on text prompts provided by the user at test time. While training these models, the text prompts usually correspond to the ground truth class names of the dataset in question. In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated. We enable Auto-Vocabulary Semantic Segmentation (AVS), in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning.
To tackle this task, we introduce the AutoSeg framework. Our framework can adapt to any Open-Vocabulary Segmentation (OVS) backbone. We propose BLIP-Cluster-Caption (BCC) to generate relevant class names, consisting of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network.
Experimental evaluations on PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] showcase the effectiveness of AutoSeg, achieving state-of-the-art results on AVS and results competitive to OVS methods which are - unlike AutoSeg- limited by the need for pre-specified class names. It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig. 2). For effective evaluation on public datasets, we furthermore propose a Large Language Model-based Auto-Vocabulary Evaluator (LAVE), tasked with mapping auto-vocabulary predictions to fixed ground truth class names in the dataset. All code will be publicly released to ensure reproducibility and to facilitate further research into AVS.
@@
In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation method for auto-vocabulary semantic segmentation which utilizes an LLM. Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for Auto-Vocabulary Semantic Segmentation on multiple public datasets.
@@ 2 Related Work
Open-Vocabulary Segmentation (OVS). OVS refers to the challenging task of segmenting images based on a list of arbitrary text queries. Contrary to traditional segmentation methods, OVS is not limited to a predefined set of classes. The field of OVS has seen significant advancements through different approaches. Approaching OVS as a zero-shot task, pioneering work by ZS3Net [1] and SPNet [33] trained custom modules that translate between visual and linguistic embedding spaces and segment by comparing word vectors and local image features. These methods attempt to learn the challenging connection between two very different and independent latent spaces. Consequently, the field of OVS has seen the introduction of methods incorporating Vision-Language Models (VLMs). VLMs pre-train a visual feature encoder with a textual feature encoder on large amounts of image-text pairs, enabling image and text representations to reside in the same feature space. One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [25], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS. LSeg [15] employed the Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments. More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [17], ZSSeg[38] and POMP [27] first obtain class-agnostic masks and use CLIP to classify each mask afterwards. Other works, including MaskCLIP [7], FC-CLIP [40], and SAN [37], focus on utilizing intermediate representations from a frozen CLIP encoder. For example, MaskCLIP uses the token features of the ViT-based CLIP encoder and compares local features to class label embeddings instead of pooling them into a global CLIP embedding. The remarkable diversity in the usage of CLIP for OVS extends to works which leverage relationships between image semantics and class labels in CLIP space [5], use CLIP for binary mask prediction [39] or ensure segmentation consistency at different granularity levels using CLIP [31]. Several works have proposed to tackle OVS outside the realm of VLMs using text-to-image diffusion models [11, 34] or custom transformer networks [35, 19, 36, 42]. Most notably, X-Decoder introduces a generalized transformer model capable of performing various vision-language tasks using a single set of parameters. One common and important limitation of the described methods is the need for textual input as a form of guidance by the user. Our work presents a novel framework which enables self-guidance by automatically finding relevant object categories using localized image captioning. Closest and concurrent to our work is the Zero-Guidance Segmentation paradigm [28], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user. Despite the similar setting, there are key differences. While we use the BLIP [16] model for both image clustering and caption generation, they design a complex pipeline consisting of image clustering with DINO, obtaining vision-language embeddings with CLIP and a custom attention-masking module, to then caption these with CLIP-guided GPT-2. This requires converting between three different latent representations throughout the process. In contrast, we achieve superior segmentation quality by enabling self-guidance using any out-of-the-box OVS model.
@@
Figure 3: Method Overview. BLIP encodings are clustered, aligned and denoised before being decoded into nouns. Generated nouns serve as self-guidance to a segmentor, which predicts the final mask. When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
@@ 3 Method
This work aims to perform auto-vocabulary semantic segmentation without any additional data, training, finetuning or textual input. To this end, we propose to identify relevant object categories in the image by captioning local image regions at different scales (Sect. 3.1), filtering the generated captions into meaningful semantic descriptions (Sect. 3.2) and using the generated classes names as target categories for a segmentor (Sect. 3.3). An overview is depicted in Fig. 3.
@@ 3.1 Local Region Captioning
To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [16]. BLIP is a powerful VLM capable of performing accurate and detailed image captioning. The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside. Subsequently, the set of embeddings is passed through transformer layers to capture contextual information across the image. Finally, all contextualized embeddings are passed to the decoder to generate text descriptions. Though effective in describing salient objects in the image, captioning models like BLIP generally fail to describe all elements of a scene comprehensively. This is likely due to the descriptions in the training data, which mostly focus on salient foreground objects. This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-ended recognition, in which many objects in the image go unnoticed. As a solution to this limitation, we propose BLIP-Cluster-Caption (BCC) to enable the BLIP text decoder to consider all encoded areas in the image. BCC clusters the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively. Finally, each BLIP feature cluster is fed to the BLIP text decoder separately. This enables generating a description of each semantically meaningful area in the image, resulting in a more comprehensive and accurate description of the image overall.
Clustering. By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$. Since common datasets often have images with higher resolution, we additionally process the image at a larger resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$. The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{H}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\hat{\textbf{B}}=\{\hat{\textbf{B}}^{\mathbf{X}_{D}},\hat{\textbf{B}}^{\mathbf{X}_{H}}\}$ at the two resolutions:
@@
$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{n}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}_{n}^{\mathbf{X}_{R}})\big{)}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{n},\quad\forall n\in\{1,2,\ldots,N\}$ (3)
where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches in resolutions $D$ and $H$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated with the function composition operator $\circ$:
$T_{l}(X)$ $=X+\textsc{MHA}(\textsc{LayerNorm}(X,X,X)$ (4)
$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}(\textsc{LayerNorm}(T_{l}(X)))$ (5)
@@
In Eq. (3), we additionally concatenate a sinusodial positional encoding [^1] Obtained with the positional-encodings Python library. [32] $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$. Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [21]:
@@
where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids. Running the clustering procedure with a 2 to 8 range for $k$ on two different image resolutions results in 14 different cluster assignments.
@@
$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (10)
@@
Cluster Denoising. To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [^2] We use the crfseg Python library. [13] and majority filter. CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them. In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches. Our implementation is specifically tailored for refining 2D segmentation map, using a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices. Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
The application of the CRF yields embeddings which are less noisy and more cohesive than the original aligned $k$-means result. To address remaining noise in the embeddings, a neighborhood majority filter is applied as a final step. For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$. The mode value from the cluster indices in that neighborhood is calculated and assigned as the new index of the central patch:
@@
Captioning. The next step involves turning clustered, denoised and enhanced BLIP embeddings into text. The BLIP text decoder is a Transformer architecture capable of processing unordered sets of embeddings of arbitrary size. We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the BLIP text decoder. Spatial information is preserved due to the presence of positional embeddings added in the clustering step. With this technique, our method essentially infers semantic categories captured by clusters and represented by BLIP embeddings. To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained. The caption generation is stochastic, with different object namings appearing in the captions depending on initialization. To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
3.2 Caption Filtering
@@
The captions generated by the BLIP text decoder are sentences in natural language. For our task, we are only interested in the class names present in each sentence. To obtain these, we filter the sentence down to relevant class names using noun extraction. Captions are parsed using the spaCy model333www.spacy.io/models/en#en_core_web_sm to provide a part-of-speech label for each word in the caption. Words labeled as nouns are kept and converted into their singular form through lemmatization. We collect all nouns generated by different clusters and cycles into one noun set and remove any duplicates, as well as nouns which do not appear in the WordNet dictionary.
3.3 Self-Guidance

### v3

@@ Abstract
Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, without training or fine-tuning. However, OVS methods typically require a human in the loop to specify the vocabulary based on the task or dataset at hand. In this paper, we introduce Auto-Vocabulary Semantic Segmentation (AVS), advancing open-ended image understanding by eliminating the necessity to predefine object categories for segmentation. Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using semantically enhanced BLIP embeddings and segments them afterwards. Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated classes and their corresponding segments. With AVS, our method sets new benchmarks on datasets PASCAL VOC, Context, ADE20K, and Cityscapes, while showing competitive performance to OVS methods that require specified class names. All code will be publicly released.
Fixed-Set Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AutoSeg (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Open-Ended Predictions Out-of-Vocabulary Predictions
Figure 1: AutoSeg Exemplary Results. AutoSeg is readily applicable to unseen images for open-ended segmentation for objects such as mascot and hole, such as the two images on the left. Furthermore, where established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment with more semantically precise object categories beyond the fixed-set ground truth, such as dachshund, bed and pagoda. Images are from the Road Anomaly [20], PASCAL [9] and ADE20K [47] datasets.
@@ 1 Introduction
While humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories, semantic segmentation methods [25] typically rely on a fixed vocabulary of predefined semantic categories. They require large human-annotated datasets and have limited capabilities for handling a broad range of classes or unknown objects. Recent studies have focused on addressing these limitations [28, 23] with models that leverage Vision-Language Models (VLMs) as an emerging category [17, 10, 7, 19, 5, 46, 3, 43, 45]. Such VLMs learn rich multi-modal features from large numbers of image-text pairs.
@@
Figure 2: Semantic Segmentation Tasks in Comparison. In traditional Semantic Segmentation, an image is segmented into fixed, predefined set of classes (fixed vocabulary). In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) should be segmented: 1) either via a human-provided prompt at runtime, or 2) the OV-method is trained to output the vocabulary of a human-annotated target dataset. In contrast, Auto-Vocabulary Segmentation automatically generates relevant object categories directly from the image. This enables true open-ended scene understanding without needing human input.
Due to the immense computational training costs of VLMs, it is common to build upon pre-trained VLMs such as CLIP [29] or BLIP [18]. However, applying these VLMs on per-pixel tasks to obtain precise locality information is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image. This makes them less useful for obtaining precise segmentation boundaries, which is crucial for downstream tasks like robot grasping. Moreover, current Open-Vocabulary segmentation (OVS) methods still require a human in the loop, inherently limiting the scalability of these methods (see Fig. 2). OVS methods can be separated into two categories: 1) prompt-based methods (e.g. [13, 16, 48]) require per-image user input to provide a known vocabulary; and 2) dataset-based methods (e.g. [10, 19, 17, 41]) require human-annotated datasets in combination with training in which the known vocabulary is baked into the method as its defined output categorization. For both categories, a human is providing a known and fixed vocabulary, which drastically simplifies the segmentation problem as the ground truth vocabulary is always provided, but it also limits the application use-cases when a human needs to be in the loop. Imagine a kitchen robot that needs to distinguish a large variety of tools and ingredients when cooking a recipe and requiring precise segmentations for grasping. It is undesirable to require a human-provided prompt for every grasp or to re-label a dataset and re-train the segmentation method every time a new recipe contains a new tool or ingredient.
To eliminate the human in the loop and fully rely on pre-trained foundation models, we propose Auto-Vocabulary Semantic Segmentation (AVS). In AVS, the unknown vocabulary is automatically generated and integrated into a semantic segmentation method thus allowing pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning. To address this task, we introduce AutoSeg, a zero-shot method that first identifies image-specific target categories and then predicts masks for them. See Fig. 1 for exemplary results. We introduce BBoost for vocabulary generation, which provides a simple yet effective semantic clustering strategy to enhance locality and semantic precision when captioning with BLIP.
[…]
@@
In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment open-ended classes in an image; 2) we propose BBoost, a novel method for generating image-specific target vocabularies, leveraging text decoding from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation approach for auto-vocabulary semantic segmentation which utilizes an LLM. Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for AVS on multiple public datasets.
Figure 3: Method Overview. BLIP encodings are clustered, aligned and denoised before being decoded into nouns by BBoost. Generated nouns serve as self-guidance to a segmentor, which predicts the final mask. When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
@@ 2 Related Work
Open-Vocabulary Segmentation (OVS). OVS tackles the challenging task of segmenting images based on arbitrary text queries rather than a predefined class set. Early OVS methods, like ZS3Net [1] and SPNet [38], treated it as a zero-shot task by training modules that align visual and linguistic embeddings, segmenting by comparing word vectors with local image features. Recently, Vision-Language Models (VLMs) pre-trained on large image-text dataset, such as Contrastive Language-Image Pre-training (CLIP)[29], have unified image and text features. CLIP has advanced the field of OVS significantly, with methods such as LSeg[17], which compares per-pixel embeddings with class embeddings, and two-stage approaches like OpenSeg [10], OPSNet [3], OVSeg [19], ZSSeg [42], and POMP [31], which first obtain class-agnostic masks and then classify each mask with CLIP. Other works, such as MaskCLIP [7], FC-CLIP [46], and SAN [43], utilize intermediate CLIP representations. Other approaches leverage CLIP to relate image semantics to class labels [5], predict binary masks [45] or ensure segmentation consistency at multiple granularities [35]. Beyond VLMs, recent models use text-to-image diffusion [12, 41] or transformers [39, 22, 40, 48], with X-Decoder providing a generalized model for various vision-language tasks.One common and significant limitation of the described methods is the need for textual input as a form of guidance by the user. Our work presents a novel framework which enables self-guidance by automatically identifying relevant object categories via localized image captioning. Closest and concurrent to our work is Zero-Guidance Segmentation [32], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user. Despite the similar setting, there are key differences: we use BLIP [18] for both image clustering and caption generation, while they use a complex pipeline involving DINO clustering, CLIP embeddings, a custom attention-masking module, and CLIP-guided GPT-2 for captioning, requiring conversions between three latent representations. In contrast, we achieve superior segmentation quality with self-guidance using any out-of-the-box OVS model.
@@ 3 Method
This work aims to perform semantic segmentation without any additional data, training, finetuning or pre-defined target categories. To this end, we propose to identify relevant object categories in an image by captioning with locality (Sect. 3.1), filter the generated captions into a meaningful vocabulary and use the vocabulary as target categories for a segmentor (Sect. 3.2). An overview is depicted in Fig. 3.
@@ 3.1 Local Region Captioning
To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [18]. BLIP is a powerful VLM capable of performing accurate and detailed image captioning. Its image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text features reside. The set of embeddings is then passed through transformer layers to capture contextual information across the image. Finally, all contextualized embeddings are decoded into text descriptions. Though effective in describing salient objects in the image, captioning models like BLIP generally fail to describe all elements of a scene comprehensively. This is likely due to the captions in the training data, which mostly focus on salient foreground objects. In turn, this limits its direct application to downstream tasks such as open-ended recognition, since many non-salient objects are missed. As a solution to this limitation, we propose BBoost for exhaustive, local captioning. BBoost clusters the encoded BLIP features into semantically meaningful feature clusters, which are enhanced to capture the object-specific feature representations more effectively. Afterwards, each individual feature cluster is fed to a pre-trained BLIP text decoder. This enables generating a description of each semantically distinct and meaningful area in the image, resulting in a more comprehensive, specific and accurate description of the image overall. Its components are detailed in the next paragraphs.
Clustering. By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$. Since common datasets often have images with higher resolution, we additionally process the image at a higher resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$. The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{H}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\hat{\textbf{B}}=\{\hat{\textbf{B}}^{\mathbf{X}_{D}},\hat{\textbf{B}}^{\mathbf{X}_{H}}\}$ at the two resolutions:
@@
$\hat{\mathbf{B}}_{n}^{\mathbf{X}_{R}}$ $=T\left(f_{\text{MLP}}\left(\mathbf{P}_{n}^{\mathbf{X}_{R}}\right)\right)\parallel\mathbf{z}_{n}^{\mathbf{X}_{R}},$
$\quad\forall R\in\{D,H\},\quad n\in\{1,\ldots,N_{R}\}$ (3)
where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the sets of patches for resolutions $D$ and $H$ respectively, and $\mathbf{P}_{n}^{\mathbf{X}_{R}}$ is the $n$-th patch. $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP), LN is LayerNorm, and $T$ is a Transformer encoder with $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated:
$T_{l}(X)$ $=X+\textsc{MHA}\big{(}\textsc{LN}(X),\textsc{LN}(X),\textsc{LN}(X)\big{)}$ (4)
$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}\big{(}\textsc{LN}(T_{l}(X))\big{)}$ (5)
@@
In Eq. (3), we concatenate, i.e. $\parallel$, a sinusoidal positional encoding [37] $\mathbf{z}_{n}^{\mathbf{X}_{R}}\in\mathbb{R}^{256}$ to each patch embedding to encode spatial information. Next, we cluster the patches in $\hat{\mathbf{B}}^{\mathbf{X}_{R}}$ for each resolution $R$ using $k$-means clustering [24]:
@@
where $C^{R}_{k}$ is the set of $k$ clusters for resolution $R$ and $\mu_{j}$ are the cluster centroids. Running the clustering procedure with $k\in\{2,\ldots,8\}$ on two different image resolutions results in 14 unique cluster assignments.
@@
$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}.$ (10)
@@
Cluster Denoising. To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [14] and majority filter. CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them. In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches. Our implementation is specifically tailored for refining 2D segmentation map, using a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices. In the pairwise potentials, we use a Gaussian filter to ensure spatial smoothness and consistency in the segmentation. The application of the CRF yields embeddings which are less noisy and more cohesive than the original aligned $k$-means result. To address remaining noise in the embeddings, a neighborhood majority filter is applied as a final step. For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$. The mode value from the cluster indices in that neighborhood is calculated and assigned as the new index of the central patch:
@@
Captioning. The next step involves turning clustered, denoised and enhanced embeddings into text. The BLIP text decoder is a transformer architecture capable of processing unordered sets of embeddings of arbitrary size. We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the text decoder. Spatial information is preserved due to the presence of positional embeddings added in the clustering step. With this technique, our method essentially infers semantic categories captured by clusters and represented by BLIP embeddings. To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning without specifically training for it. The caption generation is stochastic, with different object namings appearing in the captions depending on initialization. To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding with multiple inference cycles.
@@
Caption Filtering. The captions generated in the previous step are sentences in natural language. For our task, we are only interested in the class names present in each sentence. To obtain these, we filter the sentence down to relevant class names by extracting all nouns using part-of-speech labels for each word in the caption. Nouns are kept and converted into their singular form through lemmatization. We collect all nouns generated by different clusters and cycles into one target vocabulary and remove any duplicates, as well as nouns which do not appear in the WordNet dictionary.
3.2 Segmentation through Self-Guidance